OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 4and1/4 inches and a width of 2and5/6 inches? A.6and3/5 inches B.7and1/12 inches C.13and1/5 inches D.14and1/6 inches

OpenStudy (welshfella):

perimeter = 2 * length + 2 * width

OpenStudy (anonymous):

so i add L+L and W+W

OpenStudy (anonymous):

8 and 2/4 + 4 and 10/6

OpenStudy (anonymous):

@welshfella @livias.random

OpenStudy (welshfella):

8 + 2/4 = 8 1/2 4 + 10/6 = 5 + 4/6 = 5 2/3 so we have 5 + 8 = 13 now you need to add 1/2 and 2/3 and add it to 13

OpenStudy (welshfella):